OpenStudy (anonymous):

Find the product of (3x − 7y)^2

OpenStudy (anonymous):

Use the formula : \[(a-b)^{2} = a ^{2} - 2ab + b ^{2} \] by substituting a = 3x b = 7y. Can you do that? :-]

OpenStudy (anonymous):

9x2 − 49y2?

OpenStudy (anonymous):

@Jas9420

OpenStudy (anonymous):

Hi! (3x+7y) (3x-7y) Difference of squares = a^2 - b^2 = (a+b) (a-b) (3x)^2 - (7y)^2 9x^2 - 49y^2 (3x-7y)^2 Perfect squares (a-b)^2= a^2 -2ab + b^2 (3x)^2 -2(3x)(7y) + (7y)^2 9x^2 - 42 xy + 49y^2 (x+5)^2 Perfect squares property (a+b)^2 = a^2 + 2ab + b^2 (x)^2 + 2(x)(5) + (5)^2 x^2 + 10 x + 25 Hope this helps!

OpenStudy (anonymous):

No @allie_bear22 You get \[(3x)^{2} - 2\times3x \times 7y + (7y)^{2}\]

OpenStudy (anonymous):

@blackprincesssen so was i right? 9x2 + 49y2

OpenStudy (anonymous):

Can you simplify it now?

OpenStudy (anonymous):

@Jas9420 i know i did the rest my self and wanted to know if my answer was right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah, yes. But you did do the whole thing, right?

OpenStudy (anonymous):

i hope i help if i did

OpenStudy (anonymous):

yea u did

OpenStudy (anonymous):

tnx u guys